A random sample of 42 taxpayers claimed an average of ​$ 9,698 in medical expenses for the year. Assume the population standard deviation for these deductions was ​$2,380. Construct confidence intervals to estimate the average deduction for the population with the levels of significance shown below. a. 2 % b. 5% c. 20 %

Please help me to solve this problem using PhStat ( NOT FORMULAS)
Answer #1

Solution:

Given:

Sample Size = n = 42

Sample mean = 8696 = X

The population standard deviation for these deductions = \sigma = $2,380.

We have to construct confidence intervals to estimate the average deduction for the population with the levels of significance shown below. a. 2 % b. 5% c. 20 %

We are given levels of significance, then corresponding confidence levels are:

a) 100 - 2 = 98%

b) 100 - 5 = 95%

c) 100 - 20 = 80%

Part a) 98% confidence interval

Use following steps in PhStat (Excel)

Book1 - Excel Page Layout Formulas Data Review View Add-ins File Home Insert MegaStat About Kadd PHStat- Data Preparation Des

Select and Enter Numbers:

About Kadd MegaStat PHStat KADD- Menu Commands Estimate for the Mean, sigma kno... X D7 A B C D % Data Population Standard De

Click on OK

Thus we get following output:

Confidence Interval Estimate for the Mean
Data
Population Standard Deviation 2380
Sample Mean 9698
Sample Size 42
Confidence Level 98%
Intermediate Calculations
Standard Error of the Mean 367.2420
Z Value -2.3263
Interval Half Width 854.3326
Confidence Interval
Interval Lower Limit 8843.67
Interval Upper Limit 10552.33

Part b) 95% confidence interval:

Estimate for the Mean, sigma kno... X % Data Population Standard Deviation: 2380 Confidence Level: 95 Input Options Sample St

Confidence Interval Estimate for the Mean
Data
Population Standard Deviation 2380
Sample Mean 9698
Sample Size 42
Confidence Level 95%
Intermediate Calculations
Standard Error of the Mean 367.2420
Z Value -1.9600
Interval Half Width 719.7810
Confidence Interval
Interval Lower Limit 8978.22
Interval Upper Limit 10417.78

Part c) 80% confidence interval:

Estimate for the Mean, sigma kno... X % Data Population Standard Deviation: 2380 Confidence Level: 80 Input Options Sample St

Confidence Interval Estimate for the Mean
Data
Population Standard Deviation 2380
Sample Mean 9698
Sample Size 42
Confidence Level 80%
Intermediate Calculations
Standard Error of the Mean 367.2420
Z Value -1.2816
Interval Half Width 470.6395
Confidence Interval
Interval Lower Limit 9227.36
Interval Upper Limit 10168.64